OpenStudy (anonymous):

What is the common ratio in the following geometric sequence? 3, 12, 48, 192, ...

OpenStudy (agent0smith):

Divide any term by the term right before it.

OpenStudy (mathstudent55):

Divide each number by the previous one. What do you get?

OpenStudy (mathstudent55):

What is 12/3 ? What is 48/12 ? That is the common ratio. Once you are told a sequence is geometric, you only need to pick one pair of consecutive numbers and divide the second one by the first one to find the common ratio.

OpenStudy (jdoe0001):

\(\large \begin{array}{lllllll} &3, &12, &48, &192, &...\\ &&\uparrow&\uparrow&\uparrow\\ &&\frac{12}{3}?&\frac{48}{12}?&\frac{192}{48}? \end{array}\)